OpenStudy (anonymous):

(12/v^2-16)-(24/v-4)=3 Please help and show steps by first finding LCD!!!!

OpenStudy (anonymous):

What is LCD?

OpenStudy (anonymous):

Lowest common denominator

OpenStudy (anonymous):

They're supposed to be fractions

OpenStudy (anonymous):

(12-16v^2)/v^2-(24-4v)/v=3 (12-16v^2)-v(24-4v)=3v^2 (12-16v^2)-v(24-4v)=3v^2 12-16v^2-24v+4v^2-3v^2=o -15v^2-24v-12=o